### v2

# Intersection numbers as mixed volumes of Newton–Okounkov bodies

Abstract.
In this paper we express any intersection number $(L_{1}\cdot\ldots\cdot L_{d})$ of ample line bundles on an irreducible projective variety by the mixed volume $V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d}))$ of their Newton–Okounkov bodies. The admissible flag $Y_{\bullet}$ of subvarieties is constructed by sections of the line bundles using Bertini’s theorem, leaving a small possibility of changing the line bundles after the flag is fixed. The proof relies on the slice formula for Newton–Okounkov bodies and on mixed volume calculations in convex geometry.
2020 Mathematics Subject Classification:
14M25, 14C17, 52A39
1. Introduction
Newton–Okounkov bodies have been introduced independently by Lazarsfeld–Mustaţă [12] and by Kaveh–Khovanskii [10] based on ideas by Okounkov [13, 14] to study line bundles in algebraic geometry by convex geometric methods. Throughout this paper, let $X$ be an irreducible projective variety of dimension $d$ defined over any algebraically closed field $K$ of arbitrary characteristic. The Newton–Okounkov body of a big line bundle $L$ on $X$ is a convex body $\Delta_{Y_{\bullet}}(L)\subseteq\mathbb{R}^{d}$ that encodes much information about $L$, for example its volume. Its construction depends on the choice of an admissible flag $Y_{\bullet}$ on $X$, that is a flag
$Y_{\bullet}\colon\qquad X=Y_{0}\supsetneq Y_{1}\supsetneq Y_{2}\supsetneq\dots\supsetneq Y_{d}=\{\text{pt.}\}$. }
of irreducible subvarieties such that the point $Y_{d}$ is regular in $Y_{k}$ for all $10\leq k\leq d-$ 1. By Jow’s theorem [9, Theorem A] and by [12, Proposition 4.1(i)], two big line bundles $L_{1}$ and $L_{2}$ are numerically equivalent if and only if $\Delta_{Y_{\bullet}}(L_{1})=\Delta_{Y_{\bullet}}(L_{2})$ for all admissible flags $Y_{\bullet}$. This raises the question [19] of how intersection numbers of line bundles can be explicitly described in terms of their Newton–Okounkov bodies. In this paper, we will answer this question in the case of ample line bundles and we will deduce a partial result in the non-ample case.
Theorem 1.1.
Let $X$ be any irreducible projective variety of dimension $d$ and $L_{1},\dots,L_{d}$ ample $\mathbb{Q}$-line bundles on $X$. There exists an admissible flag $Y_{\bullet}$ on $X$ such that
$(L_{1}\cdot\ldots\cdot L_{d})=d!V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d})),$
where $(L_{1}\cdot\ldots\cdot L_{d})$ denotes the intersection number of the line bundles $L_{1},\dots,L_{d}$ and $V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d}))$ denotes the mixed volume of $\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d})$.
The theorem is a special case of the following more general and more explicit result.
Theorem 1.2.
Let $X$ be any irreducible projective variety of dimension $d$ and $L_{1},\dots,L_{k},L$ ample $\mathbb{Q}$-line bundles on $X$ with $10\leq k\leq d-$ 1. There exists an admissible flag $Y_{\bullet}$ on $X$ with the following property: For any $\mathbb{Q}$-line bundle $M$ on $X$ and any
$L_{k+1},\dots,L_{d}\in C_{L}(M)=\{\lambda L+\mu M~{}|~{}\lambda\in\mathbb{R},\mu\in\mathbb{R}_{\geq 0}\}\cap\mathrm{Amp}(X),$
where $\mathrm{Amp}(X)\subseteq N^{1}(X)_{\mathbb{R}}$ denotes the ample cone, it holds
(1.1) $(L_{1}\cdot\ldots\cdot L_{d})=d!V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d})).$
Moreover, even if $L$ is not ample, every admissible flag with the properties
  1. (i)
    $Y_{j}\subseteq Y_{j-1}$ is a Cartier divisor for every $11\leq j\leq d-$ 1.
  2. (ii)
    For every $1\leq j\leq k$ there exists an $r_{j}\in\mathbb{Q}_{>0}$ such that $r_{j}\mathcal{O}_{Y_{j-1}}(Y_{j})\equiv L_{j}|_{Y_{j-1}}$, where $\equiv$ denotes numerical equivalence of $\mathbb{Q}$-line bundles.
  3. (iii)
    For every $k+1\leq j\leq d-1$ there is an $r_{j}\in\mathbb{Q}$ such that $r_{j}\mathcal{O}_{Y_{j-1}}(Y_{j})\equiv L|_{Y_{j-1}}$.
satisfies Equation (1.1) for all $L_{k+1},\dots,L_{d}\in C_{L}(M)$ for any $\mathbb{Q}$-line bundle $M$.
To prove the theorem, we will first give the construction of a flag satisfying properties (i)–(iii) by Bertini’s theorem. Then we will show Equation (1.1) by induction on $k$, making use of the slice formula for Newton–Okounkov bodies. We already proved the case $k=0$ in [17]. We will also need a lemma about the compatibility of mixed volumes of certain convex bodies with restriction to the hyperplane $\{x_{1}=0\}$, which we will prove beforehand in Section 3.
One may ask whether there is also a version of Theorem 1.1 for big line bundles. In general, the volume $V(L)=d!V(\Delta_{Y_{\bullet}}(L))$ of a big line bundle $L$ corresponds to the volume of its Newton–Okounkov body but not to its self-intersection number $(L^{d})$. Thus, we have to replace intersection numbers by so-called positive intersection numbers. Positive intersection products have been introduced in [2] in the analytic context and in [3] in the algebraic context. Since we are only interested in their degree, we can define the positive intersection number $\langle L_{1},\dots,L_{d}\rangle$ of big line bundles $L_{1},\dots,L_{d}$ on $X$ by
$\langle L_{1}\cdot\ldots\cdot L_{d}\rangle=\sup_{(\pi,\widetilde{L}_{1},\dots,\widetilde{L}_{d})}(\widetilde{L}_{1}\cdot\ldots\cdot\widetilde{L}_{d})$
where the supremum runs over all tuples $(\pi,\widetilde{L}_{1},\dots,\widetilde{L}_{d})$, where $\pi\colon\widetilde{X}\to X$ denotes a birational morphism for a smooth projective variety $\widetilde{X}$ and $\widetilde{L}_{i}$ is a nef line bundle on $\widetilde{X}$ such that $\pi^{*}L_{i}-\widetilde{L}_{i}$ is effective for all $1\leq i\leq d$. This coincides with the classical intersection number if $L_{1},\dots,L_{d}$ are nef.
On the other hand, we will also consider a more general definition of Newton–Okounkov bodies by associating a convex body $\Delta_{\nu}(L)$ to any big line bundle $L$ on $X$ and any rank $d$ valuation $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$ instead of an admissible flag. Here, $K(X)$ denotes the function field of $X$ and $\mathbb{Z}^{d}$ is lexicographically ordered. One may ask, whether
$\langle L_{1}\cdot\ldots\cdot L_{d}\rangle=d!\sup_{\nu}V(\Delta_{\nu}(L_{1}),\dots,\Delta_{\nu}(L_{d}))$
for all big line bundles $L_{1},\dots,L_{d}$ on $X$, where the supremum runs over all rank $d$ valuations $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$. We refer to [18, Conjecture 8.8] for a more general conjecture by Xia. As a consequence of Theorem 1.1, we will show that the right-hand side is always at least as large as the left-hand side.
Corollary 1.3.
Let $X$ be any irreducible projective variety of dimension $d$. It holds
$\langle L_{1}\cdot\ldots\cdot L_{d}\rangle\leq d!\sup_{\nu}V(\Delta_{\nu}(L_{1}),\dots,\Delta_{\nu}(L_{d}))$
for all big $\mathbb{Q}$-line bundles $L_{1},\dots,L_{d}$ on $X$.
One of the main motivations of Theorem 1.1 is to translate general inequalities between mixed volumes of convex bodies to inequalities between intersection numbers of nef line bundles. As a first use case, we will derive the following corollary.
Corollary 1.4.
Let $X$ be any irreducible projective variety of dimension $d$ and $L,L_{1},\dots L_{d}$ any nef $\mathbb{R}$-line bundles on $X$. It holds
$(L_{1}\cdot\ldots\cdot L_{d})\cdot(L^{d})\leq d\cdot(L^{d-1}\cdot L_{d})\cdot(L_{1}\cdot\ldots\cdot L_{d-1}\cdot L).$
We remark that a more general inequality has been proven by Jiang–Li [8, Theorem 3.5] by different methods using multipoint Okounkov bodies.
2. Newton–Okounkov Bodies
In this section we review the definition and main properties of Newton–Okounkov bodies. We refer to [12] for more details. Moreover, we will discuss Newton–Okounkov bodies associated to valuations instead of admissible flags. We refer to [1] and [4] for more details on this part.
Let $Y_{\bullet}$ be an admissible flag on $X$ and $L$ a line bundle on $X$. There is a valuation-like function
$\nu_{Y_{\bullet}}\colon H^{0}(X,L)\setminus\{0\}\to\mathbb{Z}^{d},\qquad s\mapsto\nu_{Y_{\bullet}}(s)=(\nu_{1}(s),\dots,\nu_{d}(s)),$
which is inductively defined by $\nu_{j}(s)=\mathrm{ord}_{Y_{j}}(s_{j-1})$, where $s_{0}=s$ and $s_{j}$ is the section of the line bundle $L|_{Y_{j}}\otimes\bigotimes_{i=1}^{j}\mathcal{O}_{Y_{i-1}}(-\nu_{i}(s)Y_{i})|_{Y_{j}}$ on $Y_{j}$ induced by $s_{j-1}$ for $j\geq 1$. The Newton–Okounkov body $\Delta_{Y_{\bullet}}(L)$ of $L$ is defined by
$\Delta_{Y_{\bullet}}(L)=\mathrm{cch}\left(\bigcup_{m\geq 1}\tfrac{1}{m}\nu_{Y_{\bullet}}\left(H^{0}(X,L^{\otimes m})\setminus\{0\}\right)\right)\subseteq\left(\mathbb{R}_{\geq 0}\right)^{d},$
where $\mathrm{cch}$ denotes closed convex hull.
Let us recall some facts about Newton–Okounkov bodies. For this purpose, we assume that $L$ is big. By [12, Theorem 2.3] we have the following formula for the volumes
(2.1) $\mathrm{vol}(\Delta_{Y_{\bullet}}(L))=\tfrac{1}{d!}\mathrm{vol}(L):=\lim_{m\to\infty}\frac{\dim H^{0}(X,L^{\otimes m})}{m^{d}}.$
Note that $\mathrm{vol}(L)=(L^{d})$ if $L$ is ample. For any integer $p>0$ it holds
(2.2) $\Delta_{Y_{\bullet}}(L^{\otimes p})=p\cdot\Delta_{Y_{\bullet}}(L)$
as shown in [12, Proposition 4.1]. Hence, the definition of Newton–Okounkov bodies canonically extends to big $\mathbb{Q}$-line bundles. Since this extension is continuous on the cone of big $\mathbb{Q}$-line bundles by [12, Theorem B], Newton–Okounkov bodies are even canonically defined for big $\mathbb{R}$-line bundles.
By sending two global sections $s_{1}\in H^{0}(X,L_{1}^{\otimes m})$ and $s_{2}\in H^{0}(X,L_{2}^{\otimes m})$ of big line bundles $L_{1},L_{2}$ to $s_{1}\otimes s_{2}\in H^{0}(X,(L_{1}\otimes L_{2})^{\otimes m})$ we get an inclusion
(2.3) $\Delta_{Y_{\bullet}}(L_{1})+\Delta_{Y_{\bullet}}(L_{2})\subseteq\Delta_{Y_{\bullet}}(L_{1}\otimes L_{2}),$
where the left hand side denotes the Minkowski sum of $\Delta_{Y_{\bullet}}(L_{1})$ and $\Delta_{Y_{\bullet}}(L_{2})$.
If $d\geq 1$, we will also consider restrictions to $Y_{1}$. First, let us denote the restricted flag
$Y_{1,\bullet}\colon\qquad Y_{1}\supsetneq Y_{2}\supsetneq\dots\supsetneq Y_{d}$
on the projective variety $Y_{1}$. We define the restricted Newton–Okounkov $\Delta_{Y_{\bullet}|Y_{1}}(L)$ in $\mathbb{R}^{d-1}$ by
$\Delta_{Y_{\bullet}|Y_{1}}(L)=\mathrm{cch}\left(\bigcup_{m\geq 1}\tfrac{1}{m}\nu_{Y_{1,\bullet}}\left(\mathrm{Im}\left(H^{0}(X,L^{\otimes m})\to H^{0}(Y_{1},L^{\otimes m}|_{Y_{1}})\right)\setminus\{0\}\right)\right).$
If $L$ is ample, then the restriction map is surjective for $m\gg 0$. Hence, we get
(2.4) $\Delta_{Y_{\bullet}|Y_{1}}(L)=\Delta_{Y_{1,\bullet}}(L|_{Y_{1}}).$
Let us recall the slice formula for Newton–Okounkov bodies from [12, Theorem 4.26]. We assume that $L$ is ample and $Y_{1}\subseteq X$ is a Cartier divisor. We write
$\mu(L;Y_{1})=\sup\{s>0~{}|~{}L-s\mathcal{O}_{X}(Y_{1})\text{ is big.}\}.$. } .
Then the slice formula for $\Delta_{Y_{\bullet}}(L)$ at $\nu_{1}=\tau$ for any real $\tau\geq 0$ with $\tau\neq\mu(L;Y_{1})$ states that
(2.5) $\Delta_{Y_{\bullet}}(L)_{\nu_{1}=\tau}=\Delta_{Y_{\bullet}|Y_{1}}(L-\tau\mathcal{O}_{X}(Y_{1}))$
considered as convex bodies in $\mathbb{R}^{d-1}\cong\{\tau\}\times\mathbb{R}^{d-1}$. Note that for $\tau>\mu(L;Y_{1})$ both sides are empty sets.
More generally, we give a construction of Newton–Okounkov bodies where we replace the admissible flag by a rank $d$ valuation, that is a map $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$ such that
  1. (i)
    $\nu(fg)=\nu(f)+\nu(g)$ for all $f,g\in K(X)^{\times}$,
  2. (ii)
    $\nu(f+g)\geq\min\{\nu(f),\nu(g)\}$ for all $f,g\in K(X)^{\times}$,
  3. (iii)
    $\nu(K(X)^{\times})=\mathbb{Z}^{d}$,
  4. (iv)
    $\nu(K^{\times})=\{0\}$,
where $\mathbb{Z}^{d}$ is lexicographically ordered. The valuation $\nu$ induces a valuation-like function on $H^{0}(X,L)$ in the following way: Let $R_{\nu}\subseteq K(X)$ denote the valuation ring associated to $\nu$. By the valuative criterion of properness there is a unique morphism $\mathrm{Spec}\,R_{\nu}\to X$, which maps $\mathrm{Spec}\,K(X)$ to the generic point of $X$. The center $\mathrm{center}(\nu)$ of $\nu$ is the image of the closed point of $\mathrm{Spec}\,R_{\nu}$ in $X$. We can choose an affine neighborhood $U$ of $\mathrm{center}(\nu)$ such that $L|_{U}\cong\mathcal{O}_{U}$. Thus, we can identify $s\in H^{0}(X,L)$ with some $f\in K(X)$ up to a factor in $R_{\nu}^{\times}$. In particular, $\nu(s)=\nu(f)$ is well defined. It also does not depend on the choice of $U$. We can now define the Newton–Okounkov body $\Delta_{\nu}(L)$ by
$\Delta_{\nu}(L)=\mathrm{cch}\left(\bigcup_{m\geq 1}\tfrac{1}{m}\nu\left(H^{0}(X,L^{\otimes m})\setminus\{0\}\right)\right)\subseteq(\mathbb{R}_{\geq 0})^{d}.$
It has been proven by Boucksom [1, Corollaire 3.9] that Equation (2.1) still holds for $\Delta_{\nu}$, that is
(2.6) $\mathrm{vol}(\Delta_{\nu}(L))=\tfrac{1}{d!}\mathrm{vol}(L).$
Note that the measure $\mu_{\nu}$ appearing in Boucksom’s Corollaire is the standard Lebesgue measure on $\mathbb{R}^{d}$ by our assumption $\nu(K(X)^{\times})=\mathbb{Z}^{d}$.
Every admissible flag induces a rank $d$ valuation
$\nu_{Y_{\bullet}}\colon K(X)^{\times}\to\mathbb{Z}^{d},\qquad f\mapsto\nu_{Y_{\bullet}}(f)=(\nu_{1}(f),\dots,\nu_{d}(f))$
by setting $\nu_{1}(f)=\mathrm{ord}_{Y_{1}}(f)$ and then inductively
$f_{i}=\frac{f_{i-1}}{g_{i}^{\nu_{i}(f_{i-1})}},\qquad\nu_{i+1}(f)=\mathrm{ord}_{Y_{i+1}}(f_{i}),$
where $g_{i}=0$ is a local equation of $Y_{i}$ on $Y_{i-1}$ in an open neighborhood of $Y_{d}$. One can check that this is compatible with the notation $\nu_{Y_{\bullet}}$ from above.
Next, we will study the compatibility of Newton–Okounkov bodies with the pullback by a birational morphism $\pi\colon\widetilde{X}\to X$ for a projective variety $\widetilde{X}$. The morphism $\pi$ induces an injection of the spaces of global sections
(2.7) $\pi^{*}\colon H^{0}(X,L)\to H^{0}\left(\widetilde{X},L\right),\qquad s\mapsto\pi^{*}s$
for every line bundle $L$ on $X$. If $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$ is a rank $d$ valuation, we can consider it as a rank $d$ valuation $\widetilde{\nu}\colon K(\widetilde{X})^{\times}\to\mathbb{Z}^{d}$ by the isomorphism $K(X)\cong K(\widetilde{X})$ induced by $\pi$. This induces two valuation-like maps
$\nu\colon H^{0}(X,L)\setminus\{0\}\to\mathbb{Z}^{d},\qquad\widetilde{\nu}\colon H^{0}\left(\widetilde{X},\pi^{*}L\right)\setminus\{0\}\to\mathbb{Z}^{d}.$
Let us check that they correspond to each other under the injection (2.7). By the uniqueness property of the valuative criterion of properness, we have
$\pi(\mathrm{center}(\widetilde{\nu}))=\mathrm{center}(\nu).$
If we choose a suitable open neighborhood $\mathrm{center}(\nu)\in U$ and a trivialization of $L|_{U}$, then $\pi^{-1}(U)$ is an open neighborhood of $\mathrm{center}(\widetilde{\nu})$ and the trivialization of $L|_{U}$ induces a trivialization of $\pi^{*}L|_{\pi^{-1}(U)}$. These trivializations allow us to consider $s\in H^{0}(X,L)\setminus\{0\}$ and $\pi^{*}s$ as elements in $K(X)^{\times}$ and $K(\widetilde{X})^{\times}$ and by construction they correspond to each other under the isomorphism $K(X)\cong K(\widetilde{X})$. Thus, we get $\nu(s)=\widetilde{\nu}(\pi^{*}s)$ as desired. Hence, the injection in (2.7) induces an inclusion of Newton–Okounkov bodies
$\Delta_{\nu}(L)\subseteq\Delta_{\widetilde{\nu}}(\pi^{*}L)$
for all big line bundles $L$ on $X$. Since the volume of a line bundle is invariant under birational transformations [11, Proposition 2.2.43], we get that
$\mathrm{vol}(\Delta_{\nu}(L))=\tfrac{1}{d!}\mathrm{vol}(L)=\tfrac{1}{d!}\mathrm{vol}(\pi^{*}L)=\mathrm{vol}(\Delta_{\widetilde{\nu}}(\pi^{*}L))$
by Equation (2.6). Note that in the book [11], the author assumes that the base field is $\mathbb{C}$, but the proof of [11, Proposition 2.2.43] is still valid for any algebraically closed field. We conclude that the above inclusion of convex bodies must already be an equality
(2.8) $\Delta_{\nu}(L)=\Delta_{\widetilde{\nu}}(\pi^{*}L).$
It has been proven in [4, Theorem 2.9] that all rank $d$ valuations $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$ are equivalent to some valuation induced by an admissible flag after some birational transformation. In particular, most of the properties of Newton–Okounkov bodies, such as (2.1), (2.2), and (2.3), also hold for the more general construction $\Delta_{\nu}(\cdot)$.
3. Mixed volumes
In this section we prove a lemma about mixed volumes that will be needed in the proof of Theorem 1.2. For details on mixed volumes of convex bodies we refer to Schneider’s book [16, Chapter 5]. For any convex bodies $K_{1},\dots,K_{n}$ in $\mathbb{R}^{d}$ the volume of the Minkowski sum $t_{1}K_{1}+\dots+t_{n}K_{n}$ is a homogeneous degree $d$ polynomial in $t_{1},\dots,t_{n}$ if $t_{i}\geq 0$ for all $i$. The mixed volumes $V(K_{i_{1}},\dots,K_{i_{d}})$ for $(i_{1},\dots,i_{d})\in\{1,\dots,n\}^{d}$ are defined by the coefficients of this polynomial, that is
(3.1) $\mathrm{vol}(t_{1}K_{1}+\dots+t_{n}K_{n})=\sum_{i_{1},\dots,i_{d}=1}^{n}t_{i_{1}}\cdots t_{i_{d}}V(K_{i_{1}},\dots,K_{i_{d}})$
for all $t_{1}\geq 0,\dots,t_{n}\geq 0$, where $V(K_{i_{1}},\dots,K_{i_{d}})$ is symmetric in $K_{i_{1}},\dots,K_{i_{d}}$. It turns out that $V(\cdot,\ldots,\cdot)$ is a non-negative and multi-linear function on the cone of convex bodies in $\mathbb{R}^{d}$ and that $V(K,\dots,K)=\mathrm{vol}(K)$. For every $\tau\in\mathbb{R}$ we denote the following hyperplane
$H_{\tau}=\{(x_{1},\dots,x_{d})\in\mathbb{R}^{d}~{}|~{}x_{1}=\tau\}.$
For any convex body $K\subseteq\mathbb{R}^{d}$ we write $K_{\tau}$ for $K\cap H_{\tau}$ considered as a convex body in $\mathbb{R}^{d-1}\cong H_{\tau}$, that means $K\cap H_{\tau}=\{\tau\}\times K_{\tau}$.
Lemma 3.1.
Let $K_{1},\dots,K_{d}$ be convex bodies in $(\mathbb{R}_{\geq 0})^{d}$ and $r\in\mathbb{R}_{>0}$. If
  1. (a)
    $K_{1,\tau}=\left(1-\tfrac{\tau}{r}\right)K_{1,0}$ for all $\tau\in[0,r)$ and
  2. (b)
    $\tfrac{\tau}{r}K_{1,0}+K_{j,\tau}\subseteq K_{j,0}$ for all $2\leq j\leq d$ and all $\tau\geq 0$,
then it holds $\frac{d}{r}V(K_{1},\dots,K_{d})=V(K_{2,0},\dots,K_{d,0})$.
Proof.
The idea is to compute the volume of $t_{1}K_{1}+\dots+t_{d}K_{d}$ slice-wise and to compare the coefficients of $t_{1}\cdots t_{d}$. By the definition of the mixed volume in Equation (3.1) we have
[…]
We can decompose the slice of the Minkowski sum by
$(t_{1}K_{1}+\dots+t_{d}K_{d})\cap H_{\tau}=\bigcup_{\tau^{\prime}\in[0,\tau]}\left((t_{1}K_{1})\cap H_{\tau^{\prime}}+(t_{2}K_{2}+\dots+t_{d}K_{d})\cap H_{\tau-\tau^{\prime}}\right).$
By assumption (a) we have
$(t_{1}K_{1})\cap H_{\tau^{\prime}}=\{\tau^{\prime}\}\times(t_{1}K_{1})_{\tau^{\prime}}=\{\tau^{\prime}\}\times t_{1}K_{1,\frac{\tau^{\prime}}{t_{1}}}=\{\tau^{\prime}\}\times\left(t_{1}-\tfrac{\tau^{\prime}}{r}\right)K_{1,0}.$
Next we show a variant of assumption (b) for the Minkowski sum $t_{2}K_{2}+\dots+t_{d}K_{d}$
$\tfrac{\tau}{r}K_{1,0}+(t_{2}K_{2}+\dots+t_{d}K_{d})_{\tau}$
$=\tfrac{\tau}{r}K_{1,0}+\bigcup_{\tau_{2}+\dots+\tau_{d}=\tau}\left(t_{2}K_{2,\frac{\tau_{2}}{t_{2}}}+\dots+t_{d}K_{d,\frac{\tau_{d}}{t_{d}}}\right)$
$=\bigcup_{\tau_{2}+\dots+\tau_{d}=\tau}\left(\tfrac{\tau}{r}K_{1,0}+t_{2}K_{2,\frac{\tau_{2}}{t_{2}}}+\dots+t_{d}K_{d,\frac{\tau_{d}}{t_{d}}}\right)$
$=\bigcup_{\tau_{2}+\dots+\tau_{d}=\tau}\left(t_{2}\left(\tfrac{\tau_{2}}{t_{2}r}K_{1,0}+K_{2,\frac{\tau_{2}}{t_{2}}}\right)+\dots+t_{d}\left(\tfrac{\tau_{d}}{t_{d}r}K_{1,0}+K_{d,\frac{\tau_{d}}{t_{d}}}\right)\right)$
$\subseteq\bigcup_{\tau_{2}+\dots+\tau_{d}=\tau}(t_{2}K_{2,0}+\dots+t_{d}K_{d,0})=t_{2}K_{2,0}+\dots+t_{d}K_{d,0},$
where all $\tau_{j}$ are non-negative. Moreover, one checks that the subset $K\subseteq(\mathbb{R}_{\geq 0})^{d}$ with the slices $K_{\tau}=\frac{\tau}{r}K_{1,0}+(t_{2}K_{2}+\dots+t_{d}K_{d})_{\tau}$ is a convex body, where the slice at $\tau=0$ contains all other slices. Thus,
(3.3) $\tfrac{\tau_{1}}{r}K_{1,0}+(t_{2}K_{2}+\dots+t_{d}K_{d})_{\tau_{1}}\subseteq\tfrac{\tau_{2}}{r}K_{1,0}+(t_{2}K_{2}+\dots+t_{d}K_{d})_{\tau_{2}}$
if $\tau_{1}\geq\tau_{2}\geq 0$.
We deduce for $\tau<rt_{1}$ that
$(t_{1}K_{1})\cap H_{\tau^{\prime}}+(t_{2}K_{2}+\dots+t_{d}K_{d})\cap H_{\tau-\tau^{\prime}}$
=$=\{\tau^{\prime}\}\times\left(t_{1}-\tfrac{\tau^{\prime}}{r}\right)K_{1,0}+\{\tau-\tau^{\prime}\}\times(t_{2}K_{2}+\dots+t_{d}K_{d})_{\tau-\tau^{\prime}}$
=$=\{\tau\}\times\left(\left(t_{1}-\tfrac{\tau^{\prime}}{r}\right)K_{1,0}+(t_{2}K_{2}+\dots+t_{d}K_{d})_{\tau-\tau^{\prime}}\right)$
$\subseteq\{\tau\}\times\left(\left(t_{1}-\tfrac{\tau}{r}\right)K_{1,0}+t_{2}K_{2,0}+\dots+t_{d}K_{d,0}\right)$
with equality if $\tau^{\prime}=\tau$. For $\tau\geq rt_{1}$ we similarly get by inclusion (3.3)
$(t_{1}K_{1})\cap H_{\tau^{\prime}}+(t_{2}K_{2}+\dots+t_{d}K_{d})\cap H_{\tau-\tau^{\prime}}$
=$=\{\tau\}\times\left(\left(t_{1}-\tfrac{\tau^{\prime}}{r}\right)K_{1,0}+(t_{2}K_{2}+\dots+t_{d}K_{d})_{\tau-\tau^{\prime}}\right)$
$\subseteq\{\tau\}\times\left((t_{2}K_{2}+\dots+t_{d}K_{d})_{\tau-t_{1}r}\right)$
with equality if $\tau^{\prime}=rt_{1}$. Note that for $\tau^{\prime}>rt_{1}$ the Minkowski sum in the first line is the empty set.
We can now apply the above observations to the computation of the integral in Equation (3.2)
$\sum_{i_{1},\dots,i_{d}=1}^{d}t_{i_{1}}\cdots t_{i_{d}}V(K_{i_{1}},\dots,K_{i_{d}})$
[…]
$=\int_{0}^{rt_{1}}\sum_{i_{1},\dots,i_{d-1}=1}^{d}\widetilde{t}_{i_{1}}\cdots\widetilde{t}_{i_{d-1}}V(K_{i_{1},0},\dots,K_{i_{d-1},0})d\tau$
[…]
where $\widetilde{t}_{1}=t_{1}-\frac{\tau}{r}$ and $\widetilde{t}_{i}=t_{i}$ for $2\leq i\leq d$. We only get a $(t_{1}\cdots t_{d})$-term in the last expression if we consider the first integral and the summands with $\{i_{1},\dots,i_{d-1}\}=\{2,\dots,d\}$. Thus,
$d!V(K_{1},\dots,K_{d})=r(d-1)!V(K_{2,0},\dots,K_{d,0}).$
Dividing by $r(d-1)!$ on both sides gives the statement of the lemma. ∎
4. Construction of the flag
In this section we show the existence of an admissible flag $Y_{\bullet}$ satisfying the properties (i)–(iii) in Theorem 1.2. For simpler notation we set
$L^{\prime}_{j}=\begin{cases}L_{j}&\text{if }1\leq j\leq k,\\ L&\text{if }k+1\leq j\leq d-1.\end{cases}$
Proceeding by induction, let us assume that we have already constructed $Y_{j}$ for some $20\leq j\leq d-$ 2. By induction we may also assume that $Y_{j}\cap X_{\mathrm{reg}}\neq\emptyset$. Since $L^{\prime}_{j+1}|_{Y_{j}}$ is ample on $Y_{j}$, there exists an $\rho_{j+1}\in\mathbb{Q}_{>0}$ such that $\rho_{j+1}L^{\prime}_{j+1}|_{Y_{j}}\cong\iota_{j}^{*}\mathcal{O}(1)$ for some embedding $\iota_{j}\colon Y_{j}\to\mathbb{P}^{N_{j}}$. Hence, we will consider $Y_{j}$ as a subvariety of $\mathbb{P}^{N_{j}}$ in the following. We use the following two versions of Bertini’s theorem: By [6, Theorem 1.1] any general hyperplane $H\subseteq\mathbb{P}^{N_{j}}$ intersects $Y_{j}$ in an irreducible Cartier divisor on $Y_{j}$. By [5, Corollary 2] we have $(H\cap Y_{j})_{\mathrm{reg}}\supseteq Y_{j,\mathrm{reg}}\cap H$ for any general hyperplane $H\subseteq\mathbb{P}^{N_{j}}$. Since every hyperplane $H\subseteq\mathbb{P}^{N_{j}}$ intersects $Y_{j}$, any general hyperplane $H\subseteq\mathbb{P}^{N_{j}}$ intersects the dense open subset $Y_{j}\cap X_{\mathrm{reg}}$ of $Y_{j}$. Hence, there is a hyperplane $H_{1}\subseteq\mathbb{P}^{N_{j}}$ such that $Y_{j+1}=H_{1}\cap Y_{j}$ is an irreducible Cartier divisor on $Y_{j}$, every regular point of $Y_{j}$ lying in $Y_{j+1}$ is also regular in $Y_{j+1}$, and $Y_{j+1}\cap X_{\mathrm{reg}}\neq\emptyset$. By construction it holds $\rho_{j+1}L^{\prime}_{j+1}|_{Y_{j}}\cong\mathcal{O}_{Y_{j}}(Y_{j+1})$. By induction we end up with an admissible flag satisfying properties (i)–(iii) in Theorem 1.2 with $r_{j}=\rho_{j}^{-1}$ after setting $Y_{d}=\{p\}$ for any point $p\in Y_{d-1}\cap X_{\mathrm{reg}}$.
5. The Newton–Okounkov bodies in this paper
The purpose of this section is to show that the Newton–Okounkov bodies in Theorem 1.2 satisfy conditions (a) and (b) of Lemma 3.1.
Lemma 5.1.
Let $L_{1},\dots,L_{d}$ be ample line bundles on $X$ and $Y_{\bullet}$ an admissible flag on $X$ satisfying the corresponding properties (i)–(iii) in Theorem 1.2 with $k\geq 1$. Then it holds
  1. (a)
    $\Delta_{Y_{\bullet}}(L_{1})_{\nu_{1}=\tau}=\left(1-\frac{\tau}{r_{1}}\right)\Delta_{Y_{\bullet}}(L_{1})_{\nu_{1}=0}$ for all $\tau\in[0,r_{1})$ and
  2. (b)
    $\frac{\tau}{r_{1}}\Delta_{Y_{\bullet}}(L_{1})_{\nu_{1}=0}+\Delta_{Y_{\bullet}}(L_{j})_{\nu_{1}=\tau}\subseteq\Delta_{Y_{\bullet}}(L_{j})_{\nu_{1}=0}$ for all $2\leq j\leq d$ and all $\tau\geq 0$.
Proof.
  1. (a)
    By the numerical equivalence $L_{1}\equiv r_{1}\mathcal{O}_{X}(Y_{1})$ we have $\mu(L_{1};Y_{1})=r_{1}$. Thus, for $\tau\in[0,r_{1})$ the slice formula (2.5) implies
    (5.1) $\Delta_{Y_{\bullet}}(L_{1})_{\nu_{1}=\tau}=\Delta_{Y_{\bullet}|Y_{1}}(L_{1}-\tau\mathcal{O}_{X}(Y_{1}))=\Delta_{Y_{1,\bullet}}\left(\left(1-\tfrac{\tau}{r_{1}}\right)L_{1}|_{Y_{1}}\right),$
    where the second equality follows by Equation (2.4) and the numerical equivalence $L_{1}-\tau\mathcal{O}_{X}(Y_{1})\equiv\left(1-\frac{\tau}{r_{1}}\right)L_{1}$. Using Equation (2.2) we deduce
    $\Delta_{Y_{1,\bullet}}\left(\left(1-\tfrac{\tau}{r_{1}}\right)L_{1}|_{Y_{1}}\right)$ $=\left(1-\tfrac{\tau}{r_{1}}\right)\Delta_{Y_{1,\bullet}}\left(L_{1}|_{Y_{1}}\right)=\left(1-\tfrac{\tau}{r_{1}}\right)\Delta_{Y_{\bullet}}\left(L_{1}\right)_{\nu_{1}=0},$
    where the second equality follows from Equation (5.1) applied to $\tau=0$. This proves the first part.
  2. (b)
    As $L_{1}\equiv r_{1}\mathcal{O}_{X}(Y_{1})$, we can compute by the slice formula (2.5), Equation (2.2) for real numbers and Equations (2.3) and (2.4)
    $\tfrac{\tau}{r_{1}}\Delta_{Y_{\bullet}}(L_{1})_{\nu_{1}=0}+\Delta_{Y_{\bullet}}(L_{j})_{\nu_{1}=\tau}$ $=\Delta_{Y_{1,\bullet}}(\tau\mathcal{O}_{X}(Y_{1})|_{Y_{1}})+\Delta_{Y_{\bullet}|Y_{1}}(L_{j}-\tau\mathcal{O}_{X}(Y_{1}))$
    $\subseteq\Delta_{Y_{1,\bullet}}(\tau\mathcal{O}_{X}(Y_{1})|_{Y_{1}})+\Delta_{Y_{1,\bullet}}((L_{j}-\tau\mathcal{O}_{X}(Y_{1}))|_{Y_{1}})$
    $\subseteq\Delta_{Y_{1,\bullet}}(L_{j}|_{Y_{1}})=\Delta_{Y_{\bullet}}(L_{j})_{\nu_{1}=0}.$
    Note for the first inclusion, that we always have $\Delta_{Y_{\bullet}|Y_{1}}(L)\subseteq\Delta_{Y_{1,\bullet}}(L|_{Y_{1}})$ for any line bundle $L$ on $X$. This completes the proof of the lemma.
6. Proof of Theorem 1.2
We will prove Theorem 1.2 by induction on $k$. By Section 4 we only have to prove the second statement. If $k=0$, we proved in [17, Theorem 1.2] that the inclusion (2.3) is indeed an equality $\Delta_{Y_{\bullet}}(N_{1})+\Delta_{Y_{\bullet}}(N_{2})=\Delta_{Y_{\bullet}}(N_{1}+N_{2})$ if $N_{1},N_{2}\in C_{L}(M)$. Thus, the theorem follows from multilinearity of the intersection number and the mixed volume, see also [17, Section 4]. Hence, we may assume $k\geq 1$. In particular, $d\geq 2$. Applying the induction hypothesis to $Y_{1}$ we get
$(L_{1}\cdot\ldots\cdot L_{d})=r_{1}(L_{2}|_{Y_{1}}\cdot\ldots\cdot L_{d}|_{Y_{1}})=r_{1}(d-1)!V(\Delta_{Y_{1,\bullet}}(L_{2}),\dots,\Delta_{Y_{1,\bullet}}(L_{d})).$
Since $L_{2},\dots,L_{d}$ are ample, we get by Equation (2.4)
$V(\Delta_{Y_{1,\bullet}}(L_{2}),\dots,\Delta_{Y_{1,\bullet}}(L_{d}))=V(\Delta_{Y_{\bullet}|Y_{1}}(L_{2}),\dots,\Delta_{Y_{\bullet}|Y_{1}}(L_{d})).$
Using the slice formula (2.5) we get
$V(\Delta_{Y_{\bullet}|Y_{1}}(L_{2}),\dots,\Delta_{Y_{\bullet}|Y_{1}}(L_{d}))=V(\Delta_{Y_{\bullet}}(L_{2})_{\nu_{1}=0},\dots,\Delta_{Y_{\bullet}}(L_{d})_{\nu_{1}=0}),$
where we consider the convex bodies in the mixed volume on the right hand side as convex bodies in $\mathbb{R}^{d-1}\cong\{0\}\times\mathbb{R}^{d-1}$. By Lemmas 3.1 and 5.1 we get
$V(\Delta_{Y_{\bullet}}(L_{2})_{\nu_{1}=0},\dots,\Delta_{Y_{\bullet}}(L_{d})_{\nu_{1}=0})=dr_{1}^{-1}V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d})).$
Thus, we conclude
$(L_{1}\cdot\ldots\cdot L_{d})=d!V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d}))$
as stated in the theorem.
7. Proof of Corollary 1.3
In this section we prove Corollary 1.3. It is enough to prove that for every $\epsilon>0$, every birational morphism $\pi\colon\widetilde{X}\to X$ for a smooth projective variety $\widetilde{X}$, and all nef line bundles $\widetilde{L}_{1},\dots,\widetilde{L}_{d}$ on $\widetilde{X}$ such that $E_{i}=\pi^{*}L_{i}-\widetilde{L}_{i}$ is effective, there is a rank $d$ valuation $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$ with
(7.1) $(\widetilde{L}_{1}\cdot\ldots\cdot\widetilde{L}_{d})\leq d!V(\Delta_{\nu}(L_{1}),\dots,\Delta_{\nu}(L_{d}))+\epsilon.$
We choose an ample line bundle $A$ on $\widetilde{X}$ and a positive rational number $\delta>0$ to be determined later. Let $\widetilde{\nu}=\nu_{Y_{\bullet}}\colon K(\widetilde{X})^{\times}\to\mathbb{Z}^{d}$ be a valuation obtained from an admissible flag $Y_{\bullet}$ on $\widetilde{X}$ associated to the ample $\mathbb{Q}$-line bundles $\widetilde{L}_{1}+\delta A,\dots,\widetilde{L}_{d}+\delta A$ on $\widetilde{X}$ as in Theorem 1.1. Using Theorem 1.1 we can now compute
(7.2) $(\widetilde{L}_{1}\cdot\ldots\cdot\widetilde{L}_{d})$ $\leq((\widetilde{L}_{1}+\delta A)\cdot\ldots\cdot(\widetilde{L}_{d}+\delta A))$
$=d!V(\Delta_{\widetilde{\nu}}(\widetilde{L}_{1}+\delta A),\dots,\Delta_{\widetilde{\nu}}(\widetilde{L}_{d}+\delta A))$
$=d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}-E_{1}+\delta A),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}-E_{d}+\delta A))$
$\leq d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}+\delta A),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}+\delta A)),$
where the last inequality follows from the inclusion
$\Delta_{\widetilde{\nu}}(\pi^{*}L_{i}-E_{i}+\delta A)+\widetilde{\nu}(e_{i})\subseteq\Delta_{\widetilde{\nu}}(\pi^{*}L_{i}+\delta A)$
for any $e_{i}\in H^{0}(X,E_{i})\setminus\{0\}$ which is induced in the same way as the inclusion (2.3). Note that the mixed volume is monotonically increasing and invariant under translation in each variable.
Using the continuity of Newton–Okounkov bodies and the monotonicity and multi-linearity of the mixed volume, we can choose $\delta>0$ such that
(7.3) $d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}+\delta A),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}+\delta A))\leq d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}))+\epsilon.$
We write $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$ for the valuation obtained from $\widetilde{\nu}$ and the isomorphism $K(X)\cong K(\widetilde{X})$ induced by $\pi$. We apply Equation (2.8) to get
(7.4) $d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}))=d!V(\Delta_{\nu}(L_{1}),\dots,\Delta_{\nu}(L_{d})).$
If we now combine the inequalities (7.2) and (7.3) and Equation (7.4), we obtain the inequality in (7.1) as desired.
8. Proof of Corollary 1.4
We prove Corollary 1.4 in this section. By continuity we may assume that $L,L_{1},\dots,L_{d}$ are ample $\mathbb{Q}$-line bundles. It has been shown by Saroglou–Soprunov–Zvavitch [15, Equation 6.1] that
$V(K_{1},\dots,K_{d})\mathrm{vol}(K)\leq dV(K,\dots,K,K_{d})V(K_{1},\dots,K_{d-1},K)$
for all convex bodies $K,K_{1},\dots,K_{d}$ in $\mathbb{R}^{d}$. If $Y_{\bullet}$ is an admissible flag on $X$ satisfying (i)–(iii) of Theorem 1.2, we get by Theorem 1.2 for $k=d-1$ and by Equation (2.1)
$(L_{1}\cdot\ldots\cdot L_{d})\cdot(L^{d})$ $=(d!)^{2}V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d}))\cdot\mathrm{vol}(\Delta_{Y_{\bullet}}(L))$
$\leq d(d!)^{2}V(\Delta_{Y_{\bullet}}(L),\dots,\Delta_{Y_{\bullet}}(L),\Delta_{Y_{\bullet}}(L_{d}))$
×$\quad\times V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d-1}),\Delta_{Y_{\bullet}}(L))$
$\leq d(L^{d-1}\cdot L_{d})\cdot(L_{1}\cdot\ldots\cdot L_{d-1}\cdot L),$
where we used for the last inequality Theorem 1.2 and the inequality
$d!\cdot V(\Delta_{Y_{\bullet}}(M_{1}),\dots,\Delta_{Y_{\bullet}}(M_{1}),\Delta_{Y_{\bullet}}(M_{2}))\leq(M_{1}^{d-1}\cdot M_{2}),$
which we proved for any admissible flag $Y_{\bullet}$ and any ample line bundles $M_{1},M_{2}$ on $X$ in [17, Lemma 6.1].
Acknowledgments
I would like to thank François Ballaÿ and Mingchen Xia for useful discussions.